OpenStudy (anonymous):

Solve (x/3) + 2 > 8.

OpenStudy (kittiwitti1):

\[\frac{ x }{ 3 } + 2 > 8\]\[subtract~two~from~both~sides:~\frac{ x }{ 3 } > 8 -2~which~becomes~\frac{ x }{ 3 } > 6.\]\[now~mulitply~the~three~by~both~sides:~3 \left(\frac{ x }{ 3 }\right) > 6 \times 3\]The answer is x > 18.

OpenStudy (zzr0ck3r):

@kittiwitti1 try `\(\text{ so it looks better }\)` in math mode. \(\text{so it looks better}\)

OpenStudy (zzr0ck3r):

;)

OpenStudy (kittiwitti1):

Oh, alright. cx

OpenStudy (zzr0ck3r):

then you can have \(x, \text{ and x}\)

OpenStudy (kittiwitti1):

Thank you ^^! (: